# LCDB 1.1: A Database Illustrating Learning Curves Are More Ill-Behaved Than Previously Thought

### v1

@@ Abstract
Sample-wise learning curves plot performance versus training set size. They are useful for studying scaling laws and speeding up hyperparameter tuning and model selection. Learning curves are often assumed to be well-behaved: monotone (i.e. improving with more data) and convex. By constructing the Learning Curves Database 1.1 (LCDB 1.1), a large-scale database with high-resolution learning curves, we show that learning curves are less often well-behaved than previously thought. Using statistically rigorous methods, we observe significant ill-behavior in approximately 14% of the learning curves, almost twice as much as in previous estimates. We also identify which learners are to blame and show that specific learners are more ill-behaved than others. Additionally, we demonstrate that different feature scalings rarely resolve ill-behavior. We evaluate the impact of ill-behavior on downstream tasks, such as learning curve fitting and model selection, and find it poses significant challenges, underscoring the relevance and potential of LCDB 1.1 as a challenging benchmark for future research.
@@ 1 Introduction
non-monotone non-convex non-monotone non-convex non-monotone convex non-monotone non-convex monotone non-convex
@@
We introduce the Learning Curves Database 1.1 (LCDB 1.1) which addresses the aforementioned limitations. We incorporate four-times more training set sizes, see Figure 1. This increases the resolution of the curves, which allows us to find more ill-behavior. We argue that, depending on how learning curves are used, data-leakage is sometimes acceptable. Therefore, we introduce two database versions, one with data-leakage and the other without, and also incorporate different feature scalings. In case a performance value is missing due to an error, we justify and document it. Besides, we include the OpenML CC-18 datasets [21], which are more carefully curated datasets.
@@ 2 Preliminaries and Related Work
Several synthetic learning problems illustrate that more data does not lead to better performance [23, 14], we find such examples in carefully curated CC-18 datasets [21] (see Table  1). Peaking is such a violation, where the error rate initially decreases as the training set size increases, then rises to a peak before decreasing again [24, 25], and is also called sample-wise double descent [5]. Peaking has been proven to occur for the Fisher classifier, and the peak effect is most severe when the training set size is equal to the dimensionality $d$ [26, 27]. Double descent describes a similar phenomenon when plotting the error versus the capacity of the model [28, 1]. Other cases of monotonicity and convexity violations include dipping and phase transitions. Loog and Duin [29] introduce the concept of dipping. The error rate initially improves, then increases without recovering, even in the limit of infinite amounts of data. A phase transition means that model performance improves abruptly, causing a distinct drop in the learning curve. Phase transitions in machine learning were studied mostly in theory [30, 1], and we are not aware of any examples on real-world datasets before this work.
@@
The LCDB 1.0 [10] includes classification learning curves of various learners on numerous datasets from the OpenML platform [31, 32]. Some curves in LCDB 1.0 are missing, resulting in actually fewer than 246 datasets, possibly due to incompatibilities with sparse matrices and long compute times. Data leakage occurs because the feature imputer was fitted on the complete data. We resolve these issues with the LCDB 1.1.
@@
Task-Set [33], LCBench [34], and BUTTER [35] are datasets containing epoch-wise curves of neural networks. They are not comparable to ours, since we study sample-wise learning curves. In deep learning, the scaling law literature focuses on much larger training sets and presents much sparse learning curves [7, 8, 9]. LCDB 1.1 instead focuses on tabular data, where many classical algorithms remain competitive with deep learning [36, 37, 38, 39], and where datasets are typically smaller. Furthermore, tabular data offer unique challenges: these datasets are rare [40], and columns are often incomparable across datasets, complicating knowledge transfer [41, 42]. Meanwhile, tabular data are crucially important for industry [43]. Insights into learning curves can help estimate how much tabular data is needed [5] which is important when data is costly.
@@
[…]
@@
Table 2 gives an overview of the main differences between LCDB 1.1 and 1.0. First, we discuss data splitting, preprocessing and we justify two the versions with and without data-leakage. We briefly discuss dataset and learner selection, and end with metrics, reproducibility, and some statistics. Some details equal to LCDB 1.0 are omitted (see Appendix A). The LCDB 1.1 is available online. [^1] https://data.4tu.nl/private_datasets/V7dDlGyQJqPc_mXUAJL1MweACKG557GQtOWIVHhYpjU
@@
LCDB 1.1 with and without dl none, min-max, standard 4 times denser 28 265 4% documented
@@
[…]
@@
In LCDB 1.0, the imputer was fitted on the whole dataset. In addition, in LCDB 1.1, we apply: no scaling (abbreviated as “noFS”), min-max scaling, or standardization of features. Because of this additional preprocessing, it is even more important to discuss how to fit the preprocessor and data-leakage. When learning curves are applied for model selection and hyperparameter tuning, the goal is to reduce computation time. We can assume the user has access to the complete dataset. Fitting the preprocessor on the whole dataset can then lead to better performance and stability, and data-leakage is acceptable. However, when trying to estimate how much data is needed, we cannot assume the user has access to all data. Thus, in this case, data-leakage is not acceptable. We therefore construct two LCDB 1.1 variants, with and without data-leakage. To prevent data-leakage, preprocessors are fitted on the train set. We compare these versions in Appendix B.
@@
In LCDB 1.0, we observe that some datasets are overly easy, resulting in flat learning curves that are already converged at the first anchor. Therefore, we include all datasets of the OpenML-CC18 benchmark [21] in LCDB 1.1, called LCDB 1.1 CC-18. This benchmark was carefully curated, filtering out datasets that are overly easy, amongst other issues. The complete LCDB 1.1, referred to as LCDB 1.1 FULL, combines CC-18 with all datasets of LCDB 1.0. The LCDB 1.1 has 28 learners, see Table 3. The dummy predicts the majority class and provides a weak baseline. One-hot features violate assumptions of Naive Bayes [55], to that end we introduce mixed Naive Bayes learners, see Appendix A for all added learners.
Table 3: The 24 learners in LCDB 1.1, their estimated amount of ill-behavior (non-monotone or non-convex), and their abbreviations.
Learners (Abbreviation) Ill-behaved Ratio Learners (Abbreviation) Ill-behaved Ratio
Decision Tree (DT) 1.5% Nearest Centroid (NC) 7.9%
Extra Tree (ET) 1.9% Complement Naive Bayes (ComplementNB) 8.3%
ensemble Gradient Boosting (ens. GB) 1.9% Passive Aggressive (PA) 9.4%
ensemble Random Forest (ens. RF) 3.0% RBF Support Vector Machine (SVM_RBF) 15.8%
Stochastic Gradient Descent Classifier (SGD) 3.4% Ridge Regression Classifier (Ridge) 17.0%
@@
We use Python, scikit-learn [56], a docker image, save all package versions and provide all the code for reproducibility. [^2] https://github.com/learning-curve-research/LCDB-1.1 We fix the seed of the learner to make them reproducible. We compute: error rate, F1, AUC, and log-loss for the validation and test sets, and we store learners’ scores or probabilistic outputs. When training fails, we record the error message and set the performance to Not-A-Number (NaN). Table 4 shows the proportions of different curve shapes (their detections are discussed in the next section). Missingness refers to NaN values and is mostly caused by Naive Bayes learners that cannot handle negative features, or many learners that cannot train on data containing only a single class. While not the main point of this work, one may wonder which learners perform best, for details see Appendix C.
@@ 4 Robustly Measuring Monotonicity, Convexity, Peaking and Dipping
We slightly adapt the procedures to detect two additional patterns: dipping and peaking. Dipping corresponds to a monotonicity violation with $j$ fixed as the last anchor $N$. Peaking is characterized by a combination of convexity and monotonicity violations: we first locate a convexity violation at $(h^{*},i^{*},j^{*})$, and then verify a monotonicity violation between $h^{*}$ and $i^{*}$ and we check for significant improvement between $i^{*}$ and $j^{*}$ (similar to violation error detection, but instead checking for improvement). If all 3 conditions are satisfied, the curve is classified as peaking.
@@ 5.1 Experimental Setup
[…]
Table 4: Ill-behavior statistics of the LCDB 1.1 variants and LCDB 1.0. Since we use a significance level of $5\%$ to detect ill-behaviors, we can expect $5\%$ false positives (in the worst-case). Therefore, only numbers larger than $5\%$ are significant, which the LCDB 1.0 barely satisfies.
@@
no FS min-max FS standardization FS no FS min-max FS standardization FS no FS (interp. )
Missing 1.9% 0.0% 8.7% 3.0% 0.4% 8.7% 11.9%
Flat 5.9% 5.8% 3.4% 8.5% 7.9% 5.3% 5.2%
Non-Monotone ($\neg$ M) 11.1% 11.2% 9.2% 10.4% 11.0% 9.5% 5.1%
Non-Convex ($\neg$ C) 10.3% 9.4% 8.4% 10.8% 9.9% 8.8% 5.7%
Ill-behaved ($\neg$ M $\cup$ $\neg$ C) 14.3% 13.5% 11.2% 14.4% 14.2% 11.8% 8.1%
Peaking 4.4% 3.3% 2.9% 4.7% 3.7% 3.2% 2.5%
Dipping 7.1% 8.5% 6.3% 7.8% 9.5% 7.2% 4.6%
An overall picture of the violations can be observed in Table 4. In this section, we only discuss the no feature scaling case (“no FS”). A substantial amount of curves is non-monotone (11.1%) and non-convex (10.3%), leading to 14.3% ill-behaved curves. This is significantly larger than the significance level $\alpha$, ruling out that these curves ill-behaviors are purely caused by noise. Note that the LCDB 1.0 barely passes this bar, underlining the need for a higher resolution database like LCDB 1.1 to detect all ill-behaviors. Peaking is responsible for 4.4% and dipping is responsible for 7.1%. The amount of flat curves is reduced for the CC-18 version compared to the FULL version as expected due to more careful curation.
@@ 5.2 DA1: How many curves are significantly ill-behaved and which learners are responsible?
We also observe peaking for LDA and the Ridge classifier. This can be expected because Ridge and LDA are closely related to Fisher [57, 4.3] [58, 4.1.5] which is known to peak. Surprisingly, KNN, Naive Bayes, and Nearest Centroid also do not always behave well. For Nearest Centroid, it was known it could dip [29], but this was never observed outside of toy settings. It can be concluded that a few learners are in fact responsible for the most ill-behaving curves. The most well-behaved learners are tree-like and ensemble learners (see also Table 3).
@@
[…]

### v2

@@ Abstract
Sample-wise learning curves plot performance versus training set size. They are useful for studying scaling laws and speeding up hyperparameter tuning and model selection. Learning curves are often assumed to be well-behaved: monotone (i.e. improving with more data) and convex. By constructing the Learning Curves Database 1.1 (LCDB 1.1), a large-scale database with high-resolution learning curves including more modern learners (CatBoost, TabNet, RealMLP, and TabPFN), we show that learning curves are less often well-behaved than previously thought. Using statistically rigorous methods, we observe significant ill-behavior in approximately 15% of the learning curves, almost twice as much as in previous estimates. We also identify which learners are to blame and show that specific learners are more ill-behaved than others. Additionally, we demonstrate that different feature scalings rarely resolve ill-behavior. We evaluate the impact of ill-behavior on downstream tasks, such as learning curve fitting and model selection, and find it poses significant challenges, underscoring the relevance and potential of LCDB 1.1 as a challenging benchmark for future research.
@@ 1 Introduction
non-monotone
non-convex
non-monotone
non-convex
non-monotone
convex
non-monotone
non-convex
monotone
non-convex
@@
We introduce the Learning Curves Database 1.1 (LCDB 1.1) which addresses the aforementioned limitations. We incorporate four-times more training set sizes, see Figure 1. This increases the resolution of the curves, which allows us to find more ill-behavior. We argue that, depending on how learning curves are used, data-leakage is sometimes acceptable. Therefore, we introduce two database versions, one with data-leakage and the other without, and also incorporate different feature scalings. In case a performance value is missing due to an error, we justify and document it. Besides, we include the OpenML CC-18 datasets [21], which are more carefully curated datasets, and some more modern tabular data learners, including boosting (CatBoost [22]), deep learning (TabNet [23], RealMLP [24]), and foundation models (TabPFN v2 [25, 26]).
@@ 2 Preliminaries and Related Work
Several synthetic learning problems illustrate that more data does not lead to better performance [28, 14], we find such examples in carefully curated CC-18 datasets [21] (see Table  1). Peaking is such a violation, where the error rate initially decreases as the training set size increases, then rises to a peak before decreasing again [29, 30], and is also called sample-wise double descent [31, 5]. Peaking has been proven to occur for the Fisher classifier, and the peak effect is most severe when the training set size is equal to the dimensionality $d$ [32, 33]. Double descent describes a similar phenomenon when plotting the error versus the capacity of the model [34, 1]. Other cases of monotonicity and convexity violations include dipping and phase transitions. Loog and Duin [35] introduce the concept of dipping. The error rate initially improves, then increases without recovering, even in the limit of infinite amounts of data. Dipping has been observed for decision trees in error rate learning curves [36, 37], and for KNN in AUC learning curves [38]. A phase transition means that model performance improves abruptly, causing a distinct drop in the learning curve. Phase transitions in machine learning were studied mostly in theory [39, 1], and we are not aware of any examples on real-world datasets before this work. Beyond classification, many of the observed learning curve irregularities, such as non-monotonicity, may also arise in regression problems [40, 41, 14] and even unsupervised learning [42], challenging the assumption that more data always helps.
@@
The LCDB 1.0 [10] includes classification learning curves of various learners on numerous datasets from the OpenML platform [43, 44, 45]. Some curves in LCDB 1.0 are missing, resulting in actually fewer than 196 datasets, possibly due to incompatibilities with sparse matrices and long compute times. Data leakage occurs because the feature imputer was fitted on the complete data. We resolve these issues with the LCDB 1.1.
@@
Task-Set [46], LCBench [47], and BUTTER [48] are datasets containing epoch-wise curves of neural networks. They are not comparable to ours, since we study sample-wise learning curves. In deep learning, the scaling law literature focuses on much larger training sets and presents much sparser learning curves [7, 8, 9]. LCDB 1.1 instead focuses on tabular data, where many classical algorithms remain competitive with deep learning [49, 50, 51, 52], and where datasets are typically smaller. Furthermore, tabular data offer unique challenges: these datasets are rare [53], and columns are often incomparable across datasets, complicating knowledge transfer [54, 55]. Meanwhile, tabular data are crucially important for industry [56]. Insights into learning curves can help estimate how much tabular data is needed [5] which is important when data is costly.
@@
Learning curves are usually modeled by parametric formulas [1, 8, 57]. Popular functions are exponential and power laws, which are motivated by the well-behaved assumption [1]. Learning curve fitting can be used to estimate the amount of data needed [5]. Mohr et al. [10] identified that parametric models with 4 parameters seem to perform best for interpolation, such as $\mathrm{POW4}$, where $\hat{C}(n)=a-b(d+n)^{-c}$. The most widespread technique is least square curve-fitting using Levenberg-Marquadt [28], more advanced techniques use Bayesian techniques and neural networks [58, 59, 60]. We investigate the effect of ill-behavior on least square curve fitting.
@@
Table 2 gives an overview of the main differences between LCDB 1.1 and 1.0. First, we discuss data splitting, preprocessing and we justify two the versions with and without data-leakage. We briefly discuss dataset and learner selection, and end with metrics, reproducibility, and some statistics. Some details equal to LCDB 1.0 are omitted (see Appendix A). The LCDB 1.1 is publicly available. [^1] LCDB 1.1 dataset: https://doi.org/10.4121/3bd18108-fad0-4e4c-affd-4341fba99306
@@
LCDB 1.1 with and without dl none, min-max, standard 4 times denser 32 265 4% documented
@@
$D\xrightarrow[\text{outer seed}~m]{\text{outer split}}\left(D_{\text{train-val}}^{(m_{o})},D_{\text{test}}^{(m_{o})}\right)\quad\text{then}\quad D_{\text{train-val}}^{(m)}\xrightarrow[n\,\text{random seed}]{\text{inner split}}\left(D_{\text{train}}^{(m_{o},m_{i})},D_{\text{val}}^{(m_{o},m_{i})}\right)$
@@
In LCDB 1.0, the imputer was fitted on the whole dataset. In LCDB 1.1, we apply: no scaling (abbreviated as “noFS”), min-max scaling, or standardization of features. Because of this additional preprocessing, it is even more important to discuss how to fit the preprocessor and data-leakage. When learning curves are applied for model selection and hyperparameter tuning, the goal is to reduce computation time. We can assume the user has access to the complete dataset. Fitting the preprocessor on the whole dataset can then lead to better performance and stability, and data-leakage is acceptable. However, when trying to estimate how much data is needed, we cannot assume the user has access to all data. Thus, in this case, data-leakage is not acceptable. We therefore construct two LCDB 1.1 variants, with and without data-leakage. To prevent data-leakage, preprocessors are fitted on the train set. We compare these versions in Appendix B.
@@
In LCDB 1.0, we observe that some datasets are overly easy, resulting in flat learning curves that are already converged at the first anchor. Therefore, we include all datasets of the OpenML-CC18 benchmark [21] in LCDB 1.1, called LCDB 1.1 CC-18. This benchmark was carefully curated, filtering out datasets that are overly easy, amongst other issues. The complete LCDB 1.1, referred to as LCDB 1.1 FULL, combines CC-18 with all datasets of LCDB 1.0.
The LCDB 1.1 has 32 learners, see Table 3. The dummy predicts the majority class and provides a weak baseline. One-hot features violate assumptions of Naive Bayes [68], to that end we introduce mixed Naive Bayes learners. Moreover, we incorporate a broader set of modern tabular learners: the boosting model CatBoost [22], deep learning models such as TabNet [23] and RealMLP [24], and the foundation model TabPFN v2 [26]. According to Erickson et al. [69], CatBoost remains a strong state-of-the-art model by default, while RealMLP achieves state-of-the-art performance after tuning and ensembling. TabNet is a popular deep learning baseline [23], and TabPFN v2 is a well-performing foundation model for tabular data [26]. See Appendix A for all added learners.
All these modern learners claim robustness to differently scaled features; therefore, we only include their no scaling (noFS) and no data-leakage variants in LCDB 1.1 FULL (except TabPFN, which does not explicitly address feature scaling) [22, 23, 24, 26]. Regarding categorical features, RealMLP and TabPFN use one-hot encoding, whereas for CatBoost and TabNet we follow their suggested practice of directly feeding categorical features into the model. For implementation details see Appendix A.
Table 3: The 32 learners in LCDB 1.1 FULL (265 OpenML datasets, no scaling version), their estimated ill-behaved (non-monotone or non-convex) ratio, and their abbreviations.
Learners (Abbreviation) Ill-behaved Learners (Abbreviation) Ill-behaved
CatBoost [22] 1.5% Complement Naive Bayes (ComplementNB) 8.3%
Decision Tree (DT) 1.5% Passive Aggressive (PA) 9.4%
TabPFN v2 [26] * 1.5% Mix Complement Naive Bayes (MixComplementNB) 10.2%
Extra Tree (ET) 1.9% Mix Multinomial Naive Bayes (MixMultinomialNB) 10.6%
ensemble Gradient Boosting (ens. GB) 1.9% RBF Support Vector Machine (SVM_RBF) 15.8%
ensemble Random Forest (ens. RF) 3.0% Ridge Regression Classifier (Ridge) 17.0%
Stochastic Gradient Descent Classifier (SGD) 3.4% Mix Gaussian Naive Bayes (MixGaussianNB) 21.5%
@@
Nearest Centroid (NC) 7.9% TabNet [23] 74.3%
    The reported results cover 210 out of 265 datasets with the maximum curves length less than 10k, due to the fact that TabPFN v2 only supports datasets with up to 10k training samples, 500 features, and 10 classes. Note that some datasets included in LCDB 1.1 were used in designing its prior.
    Some LCDB 1.1 datasets were used in the meta-train benchmark for designing and meta-tuning RealMLP.
@@
We use Python, scikit-learn [70], a docker image, save all package versions and provide all the code for reproducibility. [^2] https://github.com/learning-curve-research/LCDB-1.1 We fix the seed of the learner to make them reproducible. We compute: error rate, F1, AUC, and log-loss for the validation and test sets, and we store learners’ scores or probabilistic outputs. When training fails, we record the error message and set the performance to Not-A-Number (NaN). Table 4 shows the proportions of different curve shapes (their detections are discussed in the next section). Missingness refers to NaN values and is mostly caused by Naive Bayes learners that cannot handle negative features. While not the main point of this work, one may wonder which learners perform best; see Appendix C.
@@ 4 Robustly Measuring Monotonicity, Convexity, Peaking and Dipping
Definition 3 (Peaking Phenomenon).
Peaking occurs if there exists a triplet of indices, $1\leq h<i<j\leq N$, such that:
$C(n_{i})>C(n_{h})\quad\text{and}\quad C(n_{i})>C(n_{j}).$ (4)
In this case, $C(n_{i})$ forms a local peak, indicating that the model’s performance temporarily degrades and subsequently recovers as more data is added.
Definition 4 (Dipping Phenomenon).
Dipping occurs if there exists an index $i$, $1\leq i<N$, such that:
$C(n_{i})<C(n_{N}).$ (5)
NN denotes the amount of anchors in a curve. This indicates a sustained degradation of model performance, with no recovery observed as more data is added.
Lastly, we describe how peaking and dipping are detected. Peaking is characterized by a combination of convexity and monotonicity violations: we first locate a convexity violation at $(h^{*},i^{*},j^{*})$, and then verify a monotonicity violation between $h^{*}$ and $i^{*}$ and we check for significant improvement between $i^{*}$ and $j^{*}$ (similar to violation error detection, but instead checking for improvement). If all 3 conditions are satisfied, the curve is classified as peaking. Dipping corresponds to a monotonicity violation with $j$ fixed as the last anchor $N$.
@@ 5.1 Experimental Setup
Both QDA and the Dummy classifier are excluded due to reproducibility issues and the lack of meaningful learning behavior, respectively. We do not conduct analyses of mixed Naive Bayes methods, as their curves are largely indistinguishable from standard Naive Bayes (Appendix C). We always focus on error rate learning curves in LCDB 1.1 CC-18, since its selection of datasets is more carefully curated. Results on LCDB 1.1 FULL are similar (see Appendix D). DA1 and DA2. A significance level of $\alpha=0.05$ is used throughout, and the curves are estimated using the validation set. Since we have 5 inner and 5 outer seeds, the learning curves are estimated from 25 repeats, which are aggregated together. To compare with LCDB 1.0, we interpolate the LCDB 1.0 curves to have the same length to ensure Bonferroni corrections are comparable. DA3. We closely follow the curve fitting methodology of LCDB 1.0 [10] and also use the validation set. We use the parametric models $\mathrm{POW4}$, $\mathrm{MMF4}$, and $\mathrm{WBL4}$ since they performed best. Flat curves are filtered because they are overly easy to fit, leading to very small MSEs. To detect them, we scale all learners’ curves to [0,1] range and classify it as flat if the maximum minus minimum value is below 0.05. DA4. We run successive halving to perform model selection, where the fidelity is determined by the anchor. Model selection is done using the validation set, and the selected model is evaluated using the test set.
An overall picture of the violations can be observed in Table 3 and 4. In this section, we only discuss the no feature scaling case (“no FS”). A substantial amount of curves is non-monotone (9.9%) and non-convex (11.5%), leading to 14.9% ill-behaved curves. This is significantly larger than the significance level $\alpha$, ruling out that these curves ill-behaviors are purely caused by noise. Note that the LCDB 1.0 barely passes this bar, underlining the need for a higher resolution database like LCDB 1.1 to detect all ill-behaviors. Peaking is responsible for 5.0% and dipping is responsible for 6.1%. The amount of flat curves is reduced for the CC-18 version compared to the FULL version as expected due to more careful curation.
Table 4: Ill-behavior statistics of the LCDB 1.1 variants and LCDB 1.0. Since we use a significance level of $5\%$ to detect ill-behaviors, we can expect $5\%$ false positives (in the worst-case). Therefore, only numbers larger than $5\%$ are significant, which the LCDB 1.0 barely satisfies. Note: “no FS” results include the statistics of 4 more modern learners.
@@ 5.2 DA1: How Many Curves Are Significantly Ill-Behaved and Which Learners Are Responsible?
no FS min-max FS standardization FS no FS min-max FS standardization FS no FS with interp.
Missing 2.1% 0.0% 8.7% 3.0% 0.0% 8.7% 11.9%
Flat 7.1% 5.8% 3.4% 9.9% 7.9% 5.3% 5.2%
Non-Monotone ($\neg$ M) 9.9% 11.2% 9.2% 9.6% 11.1% 9.5% 5.1%
Non-Convex ($\neg$ C) 11.5% 9.4% 8.4% 12.3% 10.0% 8.8% 5.7%
Ill-behaved ($\neg$ M $\cup$ $\neg$ C) 14.9% 13.5% 11.2% 15.4% 14.3% 11.8% 8.1%
Peaking 5.0% 3.3% 2.9% 5.7% 3.7% 3.2% 2.5%
Dipping 6.1% 8.5% 6.3% 6.9% 9.6% 7.2% 4.6%
@@
We also observe peaking for LDA and the Ridge classifier. This can be expected because Ridge and LDA are closely related to Fisher [71, 4.3] [72, 4.1.5] which is known to peak. Surprisingly, KNN, Naive Bayes, and Nearest Centroid also do not always behave well. For Nearest Centroid, it was known it could dip [35], but this was never observed outside of toy settings. It can be concluded that a few learners are in fact responsible for the most ill-behaving curves. The most well-behaved learners are tree-like and ensemble learners (see also Table 3).
While many modern learners, such as CatBoost, RealMLP, and TabPFN v2, tend to exhibit well-behaved learning curves, this is not always the case. The notably ill-behaved learner TabNet exhibits substantial non-convexity in its learning curves (see Appendix D), which mostly appears to stem from phase transition phenomena (see Appendix E.2). We suspect that TabNet was designed for large datasets, and that its default hyperparameters are not suitable for small dataset sizes. Indeed, we find that generally for large training set sizes, TabNet performs well (see Appendix C).
@@
Figure 4: Estimated probability ($\%$) of different ill-behaviors, (a) Monotonicity Violation, (b) Convexity Violation, (c) Dipping, (d) Peaking, for learners with different feature scalings. For all results see Appendix D. Observe that feature scaling for most learners does not lead to significant changes. Ridge and MLP improve significantly, while NC becomes more ill-behaved.